OpenStudy (candyme):

1 6t + 62s = --- (3t  42s), for t 2 I have done some like this, but this one is confusing me. I want to actually see HOW to do it, not just the answer.

OpenStudy (anonymous):

\[6y+62s=\frac{1}{2}(3t+42s)\]?

OpenStudy (candyme):

Yes, but the 6y is actually 6t.

OpenStudy (anonymous):

yes of course. i was just wondering about the plus sign

OpenStudy (anonymous):

wanna get rid of the fraction first? multiply by 2 to get \[12t+124s=3t+42s\]

OpenStudy (anonymous):

\[12t-3t=42x=124x\] \[11t=-81s\] \[t=\frac{-81s}{11}\] check my algebra

OpenStudy (anonymous):

yikes what a bunch of typos let me try again

OpenStudy (anonymous):

9t = 166 t = 166/9

OpenStudy (anonymous):

\[12t+124s=3t+42s\] \[12t-3t=42s-124s\] \[11t=-82s\] \[t=\frac{-82s}{11}\]

OpenStudy (anonymous):

lol i followed the typos lol...

OpenStudy (anonymous):

btw its 9t not 11 t

OpenStudy (anonymous):

yeah i suppose it is.

OpenStudy (candyme):

Ok thanks guys :)